OpenStudy (anonymous):

The roots of an equation are x = -1 ± i.

OpenStudy (anonymous):

x^2+( )+2=0

OpenStudy (anonymous):

...

OpenStudy (trojanpoem):

assume m = -1 + i , L = -1 - i the sum of them = -1 + i -1 - i = -2 the mul = (-1 + i ) ( -1 - i ) = 1 - (i^2) = 2 so it's 2